# Iterative LLM-Based Generation and Refinement of Distracting Conditions in Math Word Problems

### v1

@@ Abstract
Mathematical reasoning serves as a crucial testbed for evaluating the intelligence of large language models (LLMs), and math word problems (MWPs) represent one of the most widely used formats. Most existing MWP datasets contain only the necessary information, while problems with distracting or excessive conditions are often overlooked. Prior studies have shown that popular LLMs experience a dramatic performance drop when such distracting conditions are introduced. However, available datasets of MWPs with distracting conditions remain limited, and most exhibit low difficulty and out-of-context expressions. These shortcomings make the distracting conditions easy to detect and disregard, thereby reducing the credibility of benchmarking on these datasets. Moreover, when distracting conditions are added, the reasoning process and answers may change, requiring intensive manual effort to check and rewrite solutions.
To address these issues, we design an iterative framework that leverages LLMs to generate distracting conditions automatically. We develop a set of prompts to revise MWPs from multiple perspectives and cognitive levels, encouraging the creation of meaningful distracting conditions as well as suggestions for further refinement. A key advantage of our framework is the preservation of shared solutions between the original and revised problems: the LLMs are explicitly guided to generate distractions that do not alter the original solution, thus eliminating the need to produce new answers. This framework is efficient and easy to deploy, substantially reducing the effort required to generate MWPs with distracting conditions while maintaining high data quality.
I Introduction
@@
Large Language Models (LLMs) have become popular tools for supporting teaching and learning [8, 7]. Among various educational tasks, solving math word problems (MWPs) is a core component of STEM education [3, 5]. A math word problem is a problem presented in a scholastic context that requires applying mathematical operations to structured numerical information to derive a solution [14]. To solve MWPs, students must understand the context, extract relevant quantities and relationships, reason from the given information, and finally infer the answer [15]. Given these reasoning requirements, MWPs have become a popular testbed for LLMs to evaluate their mathematical and logical reasoning abilities. Recent studies [17, 20, 8] have focused on leveraging LLMs’ capabilities in mathematical operations and logical reasoning, benchmarking and improving their performance on such tasks. However, most existing MWPs contain only the information necessary for solving the problems, while those with irrelevant or distracting conditions remain scarce. Educational and psychological studies have shown that MWPs with distracting or irrelevant information require more advanced cognitive functions to solve [6, 11], posing additional challenges for mathematical training.
@@
Although LLMs demonstrate impressive performance in solving and explaining MWPs from elementary to undergraduate levels, irrelevant conditions remain a critical challenge for their reasoning. When irrelevant information is present, LLMs tend to incorporate it into their reasoning processes, often leading to redundant or even incorrect solutions. GSM-Symbolic [9] evaluates state-of-the-art LLMs on MWP datasets both with and without irrelevant conditions, revealing significant performance drops across models of different scales. GSM-8K-Adv [1] introduces a pipeline that generates irrelevant conditions for MWPs via adversarial augmentation, forming a loop that alternates between generation and solving, and trains a fine-tuned model on the resulting adversarial instances. GSM-IC [12] adopts a template-based approach to produce irrelevant conditions, where pre-selected topics (e.g., age or height of a human) are converted into templates and inserted into existing MWPs. Although these works generate readable and fluent conditions, the resulting problems are often of low quality. As illustrated in Fig. 1, GSM-8K-Adv includes examples with entirely off-topic conditions (e.g., humidity level of house), which are unrelated to any entities in the original problem. Since other relevant conditions focus on the same topic (number of cats), the added condition is obviously irrelevant and easily ignored. Similarly, in GSM-IC, although the question and conditions concern ages, the added irrelevant condition introduces a new character (Claire’s father). It is evident that the question is about Jessica, and only Claire—not her father—is directly related to the target relationships. Furthermore, most existing works rely on a predefined set of templates in which entities and attributes are chosen based on the creators’ intuition. This not only leads to off-topic content, as observed in GSM-8K-Adv, but also makes the datasets dull and rigid: once students or models recognize the patterns (e.g., sentences about environmental humidity are always irrelevant), the intended challenge of such problems diminishes.
Examples of MWPs with Distracting Conditions Example from GSM-8K-Adv [1] A pet store had 38 siamese cats and 25 house cats, with a humidity level of 50%. During a sale they sold 45 cats, and the humidity level rose to 70%.<<Question>> How many cats do they have left? Example from IC-GSM [12] Jessica is six years older than Claire. In two years, Claire will be 20 years old. Twenty years ago, the age of Claire’s father is 3 times of Jessica’s age. <<Question>> How old is Jessica now? Example from IGC-MWP (ours): Albert is wondering how much pizza he can eat in one day. He buys 2 large pizzas and 2 small pizzas. A large pizza has 16 slices, a small pizza has 8 slices, and a medium pizza has 12 slices. If he eats all the large and small pizzas, how many pieces does he eat that day?
@@
Figure 1: Example of MWPs with Distracting (Irrelevant) Conditions. The added irrelevant conditions are marked with underlines.
@@
II Related Work
Large language models (LLMs) have achieved strong performance in mathematical reasoning, especially with effective prompting strategies. The Chain-of-Thought (CoT) approach [17] instructs LLMs to reason step by step, improving accuracy across reasoning tasks. Least-to-Most (LTM) prompting [20] decomposes complex problems into simpler subproblems, while self-consistency [16] aggregates multiple reasoning paths through majority voting. ReAct [18] further combines reasoning with external tool use to enhance accuracy.
Beyond prompting, several studies enhance LLMs through fine-tuning. WizardMath [8] introduces Evol-Instruct, using reinforcement learning with both upward and downward evolution to expand mathematical capability. MAmmoTH [19] builds MathInstruct, an instruction-tuning dataset that jointly trains LLMs on textual and code-based reasoning. Despite these advances, most evaluations focus on MWPs with only essential information, neglecting irrelevant or distracting conditions.
Recent works reveal that LLM performance drops sharply when such distractions are added. GSM-Symbolic [9] generates MWP variants through symbolic templates; GSM-8K-Adv [1] adopts an adversarial framework to create irrelevant conditions; and GSM-IC [12] adds template-based irrelevant sentences. While efficient, these methods depend heavily on fixed templates and thus produce limited and repetitive variations.
In contrast, our framework removes topic or entity constraints and instead guides LLMs with general rubrics. This design allows more flexible and context-aware generation of distracting conditions, improving both diversity and realism.
III Framework of IGC-MWP
III-A Method Overview
We design a novel framework, IGC-MWP, for generating distracting conditions in math word problems (MWPs). The framework leverages the reasoning and generation capabilities of large language models (LLMs) to reduce intensive manual labor. All components in this pipeline are implemented using LLMs. Specifically, the model is prompted to add a distracting condition to a given MWP, and then to verify the correctness and quality of the resulting problem. Using the reasoning explanations produced during these checking steps, the LLM further improves the generated problems by identifying deficiencies and regenerating the distracting conditions as needed. The overall framework is illustrated in Fig. 2.
@@ III-A Method Overview
A key challenge in problem generation is the heavy demand for data annotation. MWPs are typically solved with paragraph-length explanations that describe the relationships among entities along with corresponding formulas or computations. To eliminate human annotation or additional solution generation, our method constrains the new MWP (with an irrelevant condition) to share the same solution as the original. This design ensures consistent solution quality while significantly reducing annotation effort.
III-B Prompt Design
@@ III-B1 Initial Generation
To further enhance task comprehension, we adopt the few-shot methodology [2, 4], incorporating one demonstration of an MWP with an added irrelevant condition in each prompt. Examples are shown in Table I.
Example Demonstration Problems
@@
Mark has a garden with flowers. He planted plants of three different colors in it. Ten of them are yellow, and there are 80% more in purple. There are only 25% as many green flowers as there are yellow and purple flowers. Additionally, the number of trees is half the number of flowers. How many flowers does Mark have in his garden?
@@
Tina makes \$18.00 an hour. If she works more than 8 hours per shift, she is eligible for overtime, paid at 1.5 times her hourly wage. If she works less than 4 hours per shift, \$2 per shift is deducted for transportation. If she works 10 hours per day for 5 days, how much money does she make?
@@
Weng earns \$12 an hour for babysitting, and earns \$5 for every 30 minutes of housecleaning. Yesterday, she only did 50 minutes of babysitting. How much did she earn?
@@
TABLE I: Demonstrations of MWPs with irrelevant conditions. Added conditions are marked by underlines.
@@
For each MWP, the original problem and its shared solution are provided to the LLM as a pair, and a revised version is generated. An example prompt is shown in Fig. 3.
@@
Initial Generation Prompts Revise this question by adding one irrelevant condition or value, while keeping the solution still the same. The irrelevant condition or value do not alter the reasoning path of the question. The irrelevant condition or value does not introduce new background stories, rather, it is related to the entities in original question. Directly give the full revised question without explanations. Example Original Question: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … How much will each of them pay if they will split the bill equally? Shared Solution: The cost of 5 pieces of hamburger is \$3 x 5 = \$<<3*5=15>>15 … So their total bill is \$15 + \$4.80 + \$2.50 +\$2.7 = <<15+4.8+2.5+2.7=25>>25. Hence, each of the five friends will contribute \$25/5 = \$<<25/5=5>>5. #### 5 Added Irrelevant Condition: There will be additional $\mathdollar 5$ service fee when the total bill is more than $\mathdollar 26$. New Question: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … There will be additional $\mathdollar 5$ service fee when the total bill is more than $\mathdollar 26$. How much will each of them pay if they will split the bill equally? Original Question: Maila is reading a 120-page book. Yesterday, she … how many pages should she read? Shared Solution: Maila read 12 x 2 = <<12*2=24>>24 pages today … Since she wants to read half of the remaining pages tomorrow, then she should read 84/2 = <<84/2=42>>42 pages. #### 42 New Question:
@@
Figure 3: Example of Step 1 Prompts (Initial Generation)
III-B2 Quantitative Relationship and Difficulty Checking
@@
After the initial generation step, we introduce two subsequent stages to evaluate and refine the revised problems at a coarse-grained level. Ideally, the added conditions should establish quantitative relationships with the entities in the original context. However, LLMs often generate trivial cases: (1) purely narrative statements without numerical relationships (e.g., "his bike is blue and gray"), or (2) quantitative but non-challenging details that simplify reasoning (e.g., "he asks for three each year" when the original problem already defines a clear pattern). Although these outputs technically satisfy the irrelevance requirement, they fail to introduce meaningful cognitive distraction.
@@
Quantitative Relationship Checking Prompts Compare the two questions, find the condition that Q2 have but Q1 does not have. Check whether this condition introduces any new quantitative relationships or new values. If so, just say YES without other explanation; if not, explain why (within 100 words). Start your response with "YES" or "NO". Q1: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … How much will each of them pay if they will split the bill equally? Q2: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … There will be additional $\mathdollar 5$ service fee when the total bill is more than $\mathdollar 26$. How much will each of them pay if they will split the bill equally?
Figure 4: Example of Step 2 Prompts: Quantitative Relationship Checking
Difficulty Level Checking Prompts These questions have the same solution; check if Q2 is more challenging to solve than Q1 by adding irrelevant conditions. Possible cases include: Q2 introduces seemingly relevant but useless conditions; Q2 introduces seemingly relevant values useless in solution; Q2 decomposes conditions or values with additional solution steps. If Q2 is more challenging than Q1, just say YES without other explanation; if not or similar, explain why (within 100 words). Start your response with "YES" or "NO". Q1: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … How much will each of them pay if they will split the bill equally? Q2: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … There will be additional $\mathdollar 5$ service fee when the total bill is more than $\mathdollar 26$. How much will each of them pay if they will split the bill equally?
Figure 5: Example of Step 3 Prompts: Difficulty Level Checking
III-B3 Desirable and Undesirable Characteristic Checking
@@
Desirable Characteristics Checking Prompts Compare the two questions, find the additional condition in Q2 compared with Q1. Check if the additional condition complies with all the following rules one by one. If it complies with all rules, just say YES; if not, explain why (within 100 words). Start your response with "YES" or "NO". Q1: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … How much will each of them pay if they will split the bill equally? Q2: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … There will be additional $\mathdollar 5$ service fee when the total bill is more than $\mathdollar 26$. How much will each of them pay if they will split the bill equally? Rules: The additional condition should be counterparts or features of existing entities in Q1; should have direct quantitative relationships with the entities in Q1; should mention something seems related to the story but useless for solving the Q2; should be consistent and reconcilable with the story in Q1; should avoid conjunctions of contrast (e.g., however, but, though); should keep the range covered by plural pronouns the same.
Figure 6: Example of Step 4 Prompts: Desirable Characteristics Checking
Undesirable Characteristics Checking Prompts Compare the two questions, find the additional condition in Q2 compared with Q1. Check if the additional condition complies with all the following rules one by one. If it complies with all rules, just say YES; if not, explain why (within 100 words). Start your response with "YES" or "NO". Q1: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … How much will each of them pay if they will split the bill equally? Q2: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … There will be additional $\mathdollar 5$ service fee when the total bill is more than $\mathdollar 26$. How much will each of them pay if they will split the bill equally? Rules: The additional condition should NOT alter the reasoning path of the original question; should NOT introduce new background stories; should NOT be impossible to lead to any other feasible calculation-needed solution even the question is changed; should NOT make the pronouns ambiguous; should NOT give explicit hints stating the additional condition is unnecessary to solve the question (e.g., ’but he doesn’t buy any’, ’but she is not interested in it’, ’but it isn’t counted in’);
Figure 7: Example of Step 5 Prompts: Undesirable Characteristics Checking
III-C Rejection Mechanism

### v3

@@ Abstract
Mathematical reasoning serves as a crucial testbed for the intelligence of large language models (LLMs), and math word problems (MWPs) are a popular type of math problems. Most MWP datasets consist of problems containing only the necessary information, while problems with distracting and excessive conditions are often overlooked. Prior works have tested popular LLMs and found a dramatic performance drop in the presence of distracting conditions. However, datasets of MWPs with distracting conditions are limited, and most suffer from lower levels of difficulty and out-of-context expressions. This makes distracting conditions easy to identify and exclude, thus reducing the credibility of benchmarking on them. Moreover, when adding distracting conditions, the reasoning and answers may also change, requiring intensive labor to check and write the solutions. To address these issues, we design an iterative framework to generate distracting conditions using LLMs. We develop a set of prompts to revise MWPs from different perspectives and cognitive levels, encouraging the generation of distracting conditions as well as suggestions for further revision. Another advantage is the shared solutions between original and revised problems: we explicitly guide the LLMs to generate distracting conditions that do not alter the original solutions, thus avoiding the need to generate new solutions. This framework is efficient and easy to deploy, reducing the overhead of generating MWPs with distracting conditions while maintaining data quality.
1 Introduction
@@
Large Language Models (LLMs) have become popular tools for supporting teaching and learning [13, 12]. Among various educational tasks, solving math word problems (MWPs) is a core component of STEM education [6, 10]. A math word problem is a problem presented in a scholastic context that requires applying mathematical operations to structured numerical information to derive a solution [23]. To solve MWPs, students must understand the context, extract relevant quantities and relationships, reason from the given information, and finally infer the answer [24]. Given these reasoning requirements, MWPs have become a popular testbed for LLMs to evaluate their mathematical and logical reasoning abilities [9, 5]. Recent studies [26, 30, 13] have focused on leveraging LLMs’ capabilities in mathematical operations and logical reasoning, benchmarking and improving their performance on such tasks. However, most existing MWPs contain only the information necessary for solving the problems, while those with irrelevant or distracting conditions remain scarce. Educational and psychological studies have shown that MWPs with distracting or irrelevant information require more advanced cognitive functions to solve [11, 17, 19], posing additional challenges for mathematical training.
@@
Although LLMs demonstrate impressive performance in solving and explaining MWPs from elementary to undergraduate levels, irrelevant conditions remain a critical challenge for their reasoning. When irrelevant information is present, LLMs tend to incorporate it into their reasoning processes, often leading to redundant or even incorrect solutions. Taking GSM-8K [9] as the base datasets, GSM-Symbolic [15] evaluates state-of-the-art LLMs on MWP datasets both with and without irrelevant conditions, revealing significant performance drops across models of different scales. GSM-8K-Adv [3] introduces a pipeline that generates irrelevant conditions for MWPs via adversarial augmentation, forming a loop that alternates between generation and solving, and trains a fine-tuned model on the resulting adversarial instances. GSM-IC [20] adopts a template-based approach to produce irrelevant conditions, where pre-selected topics (e.g., age or height of a human) are converted into templates and inserted into existing MWPs. Although these works generate readable and fluent conditions, the resulting problems are often of low quality. As illustrated in Table 1, GSM-8K-Adv includes examples with entirely off-topic conditions (e.g., humidity level of house), which are unrelated to any entities in the original problem. Since other relevant conditions focus on the same topic (number of cats), the added condition is obviously irrelevant and easily ignored. Similarly, in GSM-IC, although the question and conditions concern ages, the added irrelevant condition introduces a new character (Claire’s father). It is evident that the question is about Jessica, and only Claire—not her father—is directly related to the target relationships. Furthermore, most existing works [15, 20] rely on a predefined set of templates in which entities and attributes are chosen based on the creators’ intuition. This not only leads to off-topic content, as observed in GSM-8K-Adv, but also makes the datasets dull and rigid: once students or models recognize the patterns (e.g., sentences about environmental humidity are always irrelevant), the intended challenge of such problems diminishes.
Dataset Example Math Word Problem (MWP)
GSM-8K-Adv [3] A pet store had 38 siamese cats and 25 house cats, with a humidity level of 50%. During a sale they sold 45 cats, and the humidity level rose to 70%. <<Question>> How many cats do they have left?
IC-GSM [20] Jessica is six years older than Claire. In two years, Claire will be 20 years old. Twenty years ago, the age of Claire’s father is 3 times of Jessica’s age. <<Question>> How old is Jessica now?
IGC-MWP (Ours) Albert is wondering how much pizza he can eat in one day. He buys 2 large pizzas and 2 small pizzas. A large pizza has 16 slices, a small pizza has 8 slices, and a medium pizza has 12 slices. If he eats all the large and small pizzas, how many pieces does he eat that day?
@@
Table 1: Examples of MWPs with distracting (irrelevant) conditions from GSM-8K-Adv [3], IC-GSM [20], and IGC-MWP (ours). The added irrelevant conditions are marked with underlines.
@@
2 Related Work
Large language models (LLMs) have demonstrated remarkable performance in mathematical reasoning, particularly when equipped with curated prompting strategies [18, 1, 14]. The Chain-of-Thought (CoT) approach [26] instructs LLMs to reason step by step, substantially improving accuracy across various reasoning tasks. Least-to-Most (LTM) prompting [30] decomposes complex problems into simpler subproblems, thereby facilitating progressive reasoning. Self-consistency [25] encourages LLMs to generate multiple reasoning paths and determine the final answer through majority voting. ReAct [28] combines reasoning with external tool usage, enabling LLMs to retrieve necessary information and enhance accuracy.
Beyond prompting, several studies have improved LLMs through fine-tuning or instruction optimization. WizardMath [13] proposes an Evol-Instruct method that performs both upward and downward evolution—creating harder and easier problems—to expand a model’s mathematical capability across difficulty levels. MAmmoTH [29] constructs MathInstruct, an instruction-tuning dataset integrating 13 popular math datasets, and jointly trains LLMs on both textual and code-based reasoning tasks. Despite these advances, most evaluations have been conducted on math word problems (MWPs) that contain only essential information [2], while MWPs with irrelevant or distracting conditions remain largely unexplored.
Recent studies have shown that such distracting conditions can significantly degrade LLM reasoning. GSM-Symbolic [15] generates variants of MWPs by first extracting symbolic templates and then refilling them with new variables. GSM-8K-Adv [3] adopts an adversarial framework that alternates between generating and solving MWPs to improve both data quality and model robustness. GSM-IC [20] creates irrelevant conditions using predefined templates—such as a person’s age, height, or shoe size—and inserts them into existing MWPs. While these methods effectively augment MWPs with additional conditions, they depend heavily on fixed human-written templates and thus exhibit limited diversity and topic flexibility.
In contrast to these template-driven approaches, our framework imposes no constraints on topics or entities. Instead, it guides LLMs through general rubrics and iterative reflection, enabling more flexible, context-aware, and creative generation of distracting conditions that better mimic realistic educational settings.
3 Framework of IGC-MWP
3.1 Pipeline Overview
We design a novel framework, IGC-MWP, for generating distracting conditions in math word problems (MWPs). The framework leverages the reasoning and generation capabilities of large language models (LLMs) to reduce intensive manual labor. All components in this pipeline are implemented using LLMs. Specifically, the model is prompted to add a distracting condition to a given MWP, and then to verify the correctness and quality of the resulting problem. Using the reasoning explanations produced during these checking steps, the LLM further improves the generated problems by identifying deficiencies and regenerating the distracting conditions as needed. The overall framework is illustrated in Figure 1.
@@ 3.1 Pipeline Overview
A key challenge in problem generation is the heavy demand for data annotation [22]. MWPs are typically solved with paragraph-length explanations that describe the relationships among entities along with corresponding formulas or computations [27]. To eliminate human annotation or additional solution generation, our method constrains the new MWP (with an irrelevant condition) to share the same solution as the original. This design ensures consistent solution quality while significantly reducing annotation effort.
3.2 Prompt Design
@@ 3.2.1 Initial Generation
To further enhance task comprehension, we adopt the few-shot methodology [4, 8], incorporating one demonstration of an MWP with an added irrelevant condition in each prompt. Examples are shown in Table 2.
Revised Problems Example Math Word Problem (MWP)
@@
Problem 1 Mark has a garden with flowers. He planted plants of three different colors in it. Ten of them are yellow, and there are 80% more in purple. There are only 25% as many green flowers as there are yellow and purple flowers. Additionally, the number of trees is half the number of flowers. How many flowers does Mark have in his garden?
@@
Problem 2 Tina makes \$18.00 an hour. If she works more than 8 hours per shift, she is eligible for overtime, paid at 1.5 times her hourly wage. If she works less than 4 hours per shift, \$2 per shift is deducted for transportation. If she works 10 hours per day for 5 days, how much money does she make?
@@
Problem 3 Weng earns \$12 an hour for babysitting, and earns \$5 for every 30 minutes of housecleaning. Yesterday, she only did 50 minutes of babysitting. How much did she earn?
@@
Table 2: Examples of MWPs with irrelevant conditions in our work. The added irrelevant conditions are marked with underlines.
@@
For each MWP, the original problem and its shared solution are provided to the LLM as a pair, and a revised version is generated. An example prompt is shown in Figure 2.
@@
Initial Generation Prompts Revise this question by adding one irrelevant condition or value, while keeping the solution still the same. The irrelevant condition or value do not alter the reasoning path of the question. The irrelevant condition or value does not introduce new background stories, rather, it is related to the entities in original question. Directly give the full revised question without explanations. Example Original Question: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each ... How much will each of them pay if they will split the bill equally? Shared Solution: The cost of 5 pieces of hamburger is \$3 x 5 = \$<<3*5=15>>15 ... So their total bill is \$15 + \$4.80 + \$2.50 +\$2.7 = <<15+4.8+2.5+2.7=25>>25. Hence, each of the five friends will contribute \$25/5 = \$<<25/5=5>>5. #### 5 Added Irrelevant Condition: There will be additional $\mathdollar 5$ service fee when the total bill is more than $\mathdollar 26$. New Question: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each ... There will be additional $\mathdollar 5$ service fee when the total bill is more than $\mathdollar 26$. How much will each of them pay if they will split the bill equally? Original Question: {Question} Shared Solution: {Solution} New Question:
@@
Figure 2: Example of Step 1 Prompts Initial Generation. The added irrelevant conditions are marked with underlines.
3.2.2 Quantitative Relationship and Difficulty Checking
@@
After the initial generation step, we introduce two subsequent stages to evaluate and refine the revised problems at a coarse-grained level. Ideally, the added conditions should establish quantitative relationships with the entities in the original context [7]. However, LLMs often generate trivial cases: (1) purely narrative statements without numerical relationships (e.g., "his bike is blue and gray"), or (2) quantitative but non-challenging details that simplify reasoning (e.g., "he asks for three each year" when the original problem already defines a clear pattern). Although these outputs technically satisfy the irrelevance requirement, they fail to introduce meaningful cognitive distraction.
@@
Quantitative Relationship Checking Prompts Compare the two questions, find the condition that Q2 have but Q1 does not have. Check whether this condition introduces any new quantitative relationships or new values. If so, just say YES without other explanation; if not, explain why (within 100 words). Start your response with "YES" or "NO". Q1: {Original Question} Q2: {Revised Question}
Figure 3: Example of Step 2 Prompts Quantitative Relationship Checking
Difficulty Level Checking Prompts These questions have the same solution; check if Q2 is more challenging to solve than Q1 by adding irrelevant conditions. Possible cases include: Q2 introduces seemingly relevant but useless conditions; Q2 introduces seemingly relevant values useless in solution; Q2 decomposes conditions or values with additional solution steps. If Q2 is more challenging than Q1, just say YES without other explanation; if not or similar, explain why (within 100 words). Start your response with "YES" or "NO". Q1: {Original Question} Q2: {Revised Question}
Figure 4: Example of Step 3 Prompts Difficulty Level Checking
3.2.3 Desirable and Undesirable Characteristic Checking
@@
Desirable Characteristics Checking Prompts Compare the two questions, find the additional condition in Q2 compared with Q1. Check if the additional condition complies with all the following rules one by one. If it complies with all rules, just say YES; if not, explain why (within 100 words). Start your response with "YES" or "NO". Q1: {Original Question} Q2: {Revised Question} Rules: The additional condition should be counterparts or features of existing entities in Q1; should have direct quantitative relationships with the entities in Q1; should mention something seems related to the story but useless for solving the Q2; should be consistent and reconcilable with the story in Q1; should avoid conjunctions of contrast (e.g., however, but, though); should keep the range covered by plural pronouns the same.
Figure 5: Example of Step 4 Prompts Desirable Characteristics Checking
Undesirable Characteristics Checking Prompts Compare the two questions, find the additional condition in Q2 compared with Q1. Check if the additional condition complies with all the following rules one by one. If it complies with all rules, just say YES; if not, explain why (within 100 words). Start your response with "YES" or "NO". Q1: {Original Question} Q2: {Revised Question} Rules: The additional condition should NOT alter the reasoning path of the original question; should NOT introduce new background stories; should NOT be impossible to lead to any other feasible calculation-needed solution even the question is changed; should NOT make the pronouns ambiguous; should NOT give explicit hints stating the additional condition is unnecessary to solve the question (e.g., ’but he doesn’t buy any’, ’but she is not interested in it’, ’but it isn’t counted in’);
Figure 6: Example of Step 5 Prompts Undesirable Characteristics Checking
3.3 Rejection Mechanism